16. X is equal to 37. Assume all sales are one-time credit sales with a probability of a collection of 96 percent. The variable cost per unit is $1.67, the sales price per unit is $4.99, and the monthly interest rate is 1+ x percent. What is the NPV of a credit sale of one extra unit?

A) $3.073 B) $3.026 C) $2.981 D) $2.936 E) $2.892
Answer #1

Monthly interest rate = 37%/12 = 0.03083333333

Expected revenue = 4.99 * 0.96 = $4.7904

Profit = 4.7904 - 1.67 = $3.1204

NPV = Profit/(1 + r)

NPV = 3.1204/(1 + 0.03083333333)

NPV = 3.027065481

Option B is correct: $3.026

Can you please upvote? Thank You :-)